if a car travels from a to b at 75mph and from b to a at 50mph what is the avrage speed of the car
Answer by nerdybill(7384)

if a car travels from a to b at 75mph and from b to a at 50mph what is the avrage speed of the car
d = distance from a-b or b-a
d/75 = time from a-b
d/50 = time from b-a
.
total distance traveled: 2d
total time: d/75 + d/50
.
distance formula:
d = rt
r = d/t
.
average speed:
notice the 'd' cancels out:
2d/(d/75 + d/50)
2d/d(1/75 + 1/50)
2/(1/75 + 1/50)
multiplying by 300/300:
600/(4 + 6)
600/(10)
60 mph